# Traffic Co-Simulation Framework Empowered by Infrastructure Camera Sensing and Reinforcement Learning

### v1

@@ IV-B2 Video Detection
YOLOv8 was also tested in place of YOLOv5 as it is a newer and more powerful version of the model. The particular pretrained checkpoint used was YOLOv8m with 25.9 million parameters. The accuracy was higher with a mean absolute error of 1.73 and root mean square error of 2.80. The distribution plot of errors is shown in Figure 4b which still shows a positive skew, but the errors are much lower both in terms of extremes and on average. However, YOLOv8 was much slower to execute with the simulation taking twice as long as real-time. The simulation was run with heavy traffic using the same trained MARL agents as before. Table IV shows that the increased accuracy did result in improved system performance for all agents except those with queue reward. For the best performing agents, the average speed increased from 6.02 m/s to 6.78 m/s, which is much closer to the 7.11 achieved in the ground truth case. Total travel time decreased significantly from 208,443 seconds to 171,526 seconds and the average waiting time came down from 77.9 seconds to 41.4 seconds. In all cases, the performance of the agents using YOLOv8 was closer to the performance from agents that used ground truth observations.
TABLE IV: Simulation performance using YOLOv8 in heavy traffic scenario
@@
(a) YOLOv5 histogram of errors
(b) YOLOv8 histogram of errors
The vehicle detection method can be fine-tuned for this specific purpose by generating training data in CARLA to train a better detector. Additionally, there are more sophisticated computer vision solutions available, for example YOLOv9 [46] and Detection Transformer [47]. Our choice was based on ease of use and fast processing within the co-simulation as there are 43 cameras. The angle and cropping of each camera can also be manually adjusted to focus only on the required region of the image. For practical realism, the location was chosen as the top of the traffic light pole, however the camera can be mounted much higher to give a better vantage. An example of this is shown in Figure 5, where the left image is from a camera at 5 meters above the ground and the right one is from 10 meters, which is able to successfully detect two additional cars.
@@
A tracking algorithm like DeepSORT [48] can be integrated with detection to boost accuracy since tracking holds across frames even if detection is momentarily lost. This would require more frequent sampling, at least 5 frames per second. In the real-world implementation this would be feasible as in such distributed control systems edge computing may be used to process the data at the intersection, but in a simulation environment, performance matters. Without computer vision the simulation runs in real-time but during heavy traffic 100 seconds of simulation time corresponded to 145 seconds of wall clock time with YOLOv5 and over 200 seconds with YOLOv8 due to the combined computation load of vehicle detection from all cameras.
IV-C Discussion: Extension of Co-simulation Framework to Digital Twin
@@
Co-simulation is a useful, and perhaps necessary, step before real-world testing for transportation research. It can enable thorough examination of automation processes that involve sensing and control in a variety of scenarios that are likely to be faced after deployment. Artificial intelligence powered applications including traffic signal control, variable speed limits [49], parking management, etc. all rely on sensing and control feedback loops integrated into the diverse urban environment. There can be many variables that affect the system and noisy sensing is one example. Robustness to those conditions should be proven in co-simulation environments for practical feasibility. The framework developed in this paper lays the foundation for further extension into the domain of digital twin technologies. A digital twin for urban traffic systems involves the real-time integration of physical infrastructure and simulation environments, enabling continuous monitoring, simulation, and optimization of the entire traffic network. Through edge computing, real-world sensors like infrastructure cameras, connected autonomous vehicles (CAVs), and even roadside units can provide continuous streams of traffic data [33]. This real-time data can be integrated with the co-simulation environment, enabling online learning and adaptive control strategies.
@@
Furthermore, connected autonomous vehicles (CAVs) could play a pivotal role in enhancing the co-simulation by acting as mobile sensors. These vehicles are equipped with advanced sensor arrays, including LIDAR, radar, and cameras, allowing them to gather rich environmental data as they navigate the traffic network [50]. By leveraging cooperative sensing from CAVs, the system can implement distributed learning algorithms that optimize traffic signal timing and traffic flow in real-time. CAVs can also be used to detect and mitigate traffic bottlenecks by actively controlling traffic patterns, especially on highways, to reduce shockwaves and smooth out traffic congestion [51].
A complete digital twin co-simulation framework in the context of traffic signal control will consist of real-time data collection, simulation, and feedback loops for continuous optimization of traffic networks. Sensing infrastructure can be enhanced with cameras and CAV data in addition to loop detectors by aggregating and preprocessing on roadside edge computing units. The edge computer will host the vehicle detection module and compare and combine results with loop detectors, traffic signal cabinets, and connected vehicle data before transmitting the relevant information to the centralized digital twin. At the central server the data from all edge devices will be organized and used for online reinforcement learning. Meanwhile the co-simulation will simultaneously be performed on the digital twin providing synthetic variation to the collected data to further aid training. In addition to MARL for TSC, the digital twin platform could enable the development of short-term traffic forecasting algorithms [52].
@@ V Conclusion
The framework described in this paper can be applied and extended to traffic signal control of real world locations and support more complex scenarios involving cooperative perception, connected autonomous vehicles, and digital twins with field collected data.

### v2

@@ 4.2.2 Video Detection
YOLOv8 was also tested in place of YOLOv5 as it is a newer and more powerful version of the model. The particular pretrained checkpoint used was YOLOv8m with 25.9 million parameters. The accuracy was higher with a mean absolute error of 1.73 and root mean square error of 2.80. The distribution plot of errors is shown in Figure 4b which still shows a positive skew, but the errors are much lower both in terms of extremes and on average. However, YOLOv8 was much slower to execute with the simulation taking twice as long as real-time. The simulation was run with heavy traffic using the same trained MARL agents as before. Table 4 shows that the increased accuracy did result in improved system performance for all agents except those with queue reward. For the best performing agents, the average speed increased from 6.02 m/s to 6.78 m/s, which is much closer to the 7.11 achieved in the ground truth case. Total travel time decreased significantly from 208,443 seconds to 171,526 seconds and the average waiting time came down from 77.9 seconds to 41.4 seconds. In all cases, the performance of the agents using YOLOv8 was closer to the performance from agents that used ground truth observations.
(a) YOLOv5 histogram of errors
(b) YOLOv8 histogram of errors
Table 4: Simulation performance using YOLOv8 in heavy traffic scenario
@@
The vehicle detection method can be fine-tuned for this specific purpose by generating training data in CARLA to train a better detector. Additionally, there are more sophisticated computer vision solutions available, for example YOLOv9 (C.-Y. Wang & Liao, 2024) and Detection Transformer (Ouyang-Zhang et al., 2022). Our choice was based on ease of use and fast processing within the co-simulation as there are 43 cameras. The angle and cropping of each camera can also be manually adjusted to focus only on the required region of the image. For practical realism, the location was chosen as the top of the traffic light pole, however the camera can be mounted much higher to give a better vantage. An example of this is shown in Figure 5, where the left image is from a camera at 5 meters above the ground and the right one is from 10 meters, which is able to successfully detect two additional cars.
@@
A tracking algorithm like DeepSORT (Wojke et al., 2017) can be integrated with detection to boost accuracy since tracking holds across frames even if detection is momentarily lost. This would require more frequent sampling, at least 5 frames per second. In the real-world implementation this would be feasible as in such distributed control systems edge computing may be used to process the data at the intersection, but in a simulation environment, performance matters. Without computer vision the simulation runs in real-time but during heavy traffic 100 seconds of simulation time corresponded to 145 seconds of wall clock time with YOLOv5 and just over 200 seconds with YOLOv8 due to the combined computation load of vehicle detection from all cameras.
4.3 Analysis of Learned Policies and Failure Modes
4.3.1 Interpretability and Learned Signal Patterns
The Q-learning algorithm employed in this study inherently supports a degree of interpretability through its Q-table structure. Each Q-table is a matrix that maps specific environmental states to the expected cumulative reward for taking a particular action from that state. The state of an agent is defined by the current phase index, a boolean indicating minimum green time elapsed, and vectors for occupancy and queue length in incoming lanes. The action space corresponds to the index of the next green phase to be activated. By inspecting the Q-table for a given intersection, one can observe the learned policy directly: for any given combination of phase, minimum green time status, and observed lane occupancy and queue lengths, the table indicates the action (next phase) that the agent has learned to be optimal. For example, an entry in the Q-table looks like this:
$\text{State: }(0,1,0,7,0,9,0,7,0,0)\rightarrow\text{Q-values: }(51.12,0.01)$
Meaning that the signal is in phase 0, corresponding to a green light for East-West traffic, the minimum green time has elapsed, and there are 7 cars coming from the West which are all stopped to form a queue, while there are 9 cars coming in from the East which are all moving as the queue value is zero. In this scenario the Q-value is determined to be high for remaining in phase zero.
This direct mapping allows for a transparent understanding of the agent’s decision-making process under various traffic conditions. For instance, if the Q-table consistently shows a high Q-value for extending a green phase when a high occupancy and queue are detected on that approach, it indicates the agent has learned to prioritize clearing that congestion. Conversely, a low Q-value for a phase might indicate the agent has learned to avoid it under certain conditions. This explicit, tabular representation of the policy contributes to the regulatability of the control policy, as changes in signal assignment can be traced back to specific changes in the observed traffic state, similar to how traditional actuated systems operate based on predefined logic. The observed robustness of agents trained with the ‘average speed’ reward function further suggests that these agents develop more stable and predictable policies, as they are less sensitive to instantaneous noise in observations compared to agents relying on more volatile metrics like waiting time or pressure.
4.3.2 Failure Modes
Phase Lock-in: Phase lock-in refers to a scenario where a traffic signal becomes stuck in a particular phase, preventing other movements from receiving a green light, leading to severe congestion. The implemented framework includes a specific error handling mechanism to mitigate this risk. As mentioned earlier, if a state is not found in an agent’s Q-table, no action is taken, but if inaction leads to a phase being stuck for over 60 seconds, phase change is forced. This rule acts as a safeguard, ensuring that even if the MARL agent fails to provide a new action or gets into an unforeseen state, the system will revert to a basic, time-based control to prevent a complete gridlock at the intersection. This mechanism directly addresses the potential for phase lock-in by enforcing a maximum green time, a common practice in traditional traffic signal control.
Queue Spillback: Queue spillback occurs when a queue of vehicles extends from one intersection back into an upstream intersection, blocking traffic flow and exacerbating congestion across the network. The study observed an instance of this failure mode with the ‘queue length based reward’ agent. Specifically, during heavy congestion, traffic backed up onto the highway, which was outside the observation range of the traffic light’s camera. Consequently, the agent consistently observed a ‘full queue’ state, and any action it took did not appear to change the queue size in its immediate observation. This led to the agent failing to learn the effective value of giving a green phase to the incoming traffic, as its reward signal was not sufficiently responsive to the actual queue dynamics extending beyond its local sensing. This highlights a critical limitation of purely local observation and reward functions in preventing network-wide congestion phenomena like spillback. While the current framework demonstrated this failure mode, future work can explore strategies to address it, such as incorporating a broader observation area, using network-level reward signals, or implementing explicit queue management strategies within the MARL framework, as some existing RL approaches for ramp metering have shown success in preventing queue spillback by estimating lower bounds for metering rates and replacing actions that violate these constraints.
4.4 Discussion: Extension of Co-simulation Framework to Digital Twin
@@
Co-simulation is a useful, and perhaps necessary, step before real-world testing for transportation research. It can enable thorough examination of automation processes that involve sensing and control in a variety of scenarios that are likely to be faced after deployment. Artificial intelligence powered applications including traffic signal control, variable speed limits (Zhang et al., 2024), parking management, etc. all rely on sensing and control feedback loops integrated into the diverse urban environment. There can be many variables that affect the system and noisy sensing is one example. Robustness to those conditions should be proven in co-simulation environments for practical feasibility. The framework developed in this paper lays the foundation for further extension into the domain of digital twin technologies. A digital twin for urban traffic systems involves the real-time integration of physical infrastructure and simulation environments, enabling continuous monitoring, simulation, and optimization of the entire traffic network. Through edge computing, real-world sensors like infrastructure cameras, connected autonomous vehicles (CAVs), and even roadside units can provide continuous streams of traffic data (Azfar et al., 2023). This real-time data can be integrated with the co-simulation environment, enabling online learning and adaptive control strategies.
@@
Such digital twin frameworks are directly applicable to CAV deployment, as they enable evaluation of CAV-infrastructure interactions under realistic sensing conditions, facilitating safer and smoother transitions from simulation to field implementation. Furthermore, CAVs could play a pivotal role in enhancing the co-simulation by acting as mobile sensors. These vehicles are equipped with advanced sensor arrays, including LIDAR, radar, and cameras, allowing them to gather rich environmental data as they navigate the traffic network (Ke, Cui, et al., 2023). By leveraging cooperative sensing from CAVs, the system can implement distributed learning algorithms that optimize traffic signal timing and traffic flow in real-time. CAVs can also be used to detect and mitigate traffic bottlenecks by actively controlling traffic patterns, especially on highways, to reduce shockwaves and smooth out traffic congestion (Zheng et al., 2020).
A complete digital twin co-simulation framework in the context of traffic signal control will consist of real-time data collection, simulation, and feedback loops for continuous optimization of traffic networks. Sensing infrastructure can be enhanced with cameras and CAV data in addition to loop detectors by aggregating and preprocessing on roadside edge computing units. The edge computer will host the vehicle detection module and compare and combine results with loop detectors, traffic signal cabinets, and connected vehicle data before transmitting the relevant information to the centralized digital twin. At the central server the data from all edge devices will be organized and used for online reinforcement learning. Meanwhile the co-simulation will simultaneously be performed on the digital twin providing synthetic variation to the collected data to further aid training. In addition to MARL for TSC, the digital twin platform could enable the development of short-term traffic forecasting algorithms (Vlahogianni et al., 2014).
@@ 5 Conclusion
Beyond advancing adaptive signal control, the findings underscore the potential of infrastructure assisted reinforcement learning to support automated vehicles. By showing that vision driven control remains effective with imperfect sensing, the study contributes a practical step toward scalable CAV deployment in real-world urban networks.
The framework described in this paper can be applied and extended to traffic signal control of real world locations and support more complex scenarios involving cooperative perception, connected autonomous vehicles, and digital twins with field collected data. This study lays the groundwork for a scalable co-simulation framework integrating vision-based sensing with MARL for traffic control. Future work will focus on enhancing the perception pipeline through post-processing techniques such as confidence-weighted filtering and multi-object tracking (e.g., DeepSORT) to improve detection accuracy and temporal consistency. Incorporating custom-trained object detectors optimized for the CARLA environment may further reduce domain mismatch. Additionally, exploring the interplay between noise characteristics, state encoding, and reward formulation, particularly under highly congested conditions, could yield new insights into designing more robust MARL agents. Real-time feasibility of these enhancements could be supported through edge computing or sensor hardware optimization.